# Looks Too Good To Be True: An Information-Theoretic Analysis of Hallucinations in Generative Restoration Models

### v2

@@ Abstract
The pursuit of high perceptual quality in image restoration has driven the development of revolutionary generative models, capable of producing results often visually indistinguishable from real data. However, as their perceptual quality continues to improve, these models also exhibit a growing tendency to generate hallucinations – realistic-looking details that do not exist in the ground truth images. The presence of hallucinations introduces uncertainty regarding the reliability of the models’ predictions, raising major concerns about their practical application. In this paper, we employ information-theory tools to investigate this phenomenon, revealing a fundamental tradeoff between uncertainty and perception. We rigorously analyze the relationship between these two factors, proving that the global minimal uncertainty in generative models grows in tandem with perception. In particular, we define the inherent uncertainty of the restoration problem and show that attaining perfect perceptual quality entails at least twice this uncertainty. Additionally, we establish a relation between mean squared-error distortion, uncertainty and perception, through which we prove the aforementioned uncertainly-perception tradeoff induces the well-known perception-distortion tradeoff. This work uncovers fundamental limitations of generative models in achieving both high perceptual quality and reliable predictions for image restoration. We demonstrate our theoretical findings through an analysis of single image super-resolution algorithms. Our work aims to raise awareness among practitioners about this inherent tradeoff, empowering them to make informed decisions and potentially prioritize safety over perceptual performance.
@@ 1 Introduction
Restoration tasks and inverse problems impact many scientific and engineering disciplines, as well as healthcare, education, communication and art. Generative artificial intelligence has transformed the field of inverse problems due to its unprecedented ability to infer missing information and restore corrupted data. In the realm of image restoration, the quest for high perceptual quality has led to a new generation of generative models  [33, 65, 18, 17, 70], capable of producing outputs of remarkable realism, virtually indistinguishable from true images.
@@
Concerns surrounding hallucinations have prompted the development of uncertainty quantification methods, designed to evaluate the reliability of generated outputs. These approaches offer crucial insights into the model’s confidence in its predictions, empowering users to assess potential deviations from the original data and make informed decisions. Despite these advancements, the intricate relationship between achieving high perceptual quality and the extent of uncertainty remains an understudied area.
@@
    We establish a relationship between uncertainty and mean squared error (MSE) distortion, demonstrating that the uncertainty-perception trade-off induces the well-known distortion-perception trade-off (Theorem 4). We empirically validate all theoretical findings through experiments with image super-resolution algorithms (Section 4.3).
@@
Figure 1: Illustration of Theorem 3. In inverse problems, the minimal attainable uncertainty exhibits a lower bound that begins at the inherent uncertainty $U_{\text{Inherent}}$ of the problem (Definition 2) and increases up to twice this value as we approach perfect perceptual quality of the recovery.
@@ 2 Related Work
Perception Quantification   Perceptual quality in restoration tasks encompasses how humans perceive the output, considering visual fidelity, similarity to the original, and absence of artifacts. While traditional metrics like PSNR and SSIM [72] capture basic similarity, they miss finer details and higher-level structures. Learned metrics like LPIPS [75], VGG-loss [64], and DISTS [23] offer improvements but still operate on pixel or patch level, potentially overlooking holistic aspects. Recently, researchers have leveraged image-level embeddings from large vision models like DINO [15] and CLIP [54] to capture high-level similarity. Here, we follow previous works [51, 12, 31] and adopt a mathematical notion of perceptual quality defined as the divergence between probability densities.
@@
The Distortion-Perception Tradeoff   The most relevant studies to our research are the work on the distortion-uncertainty tradeoff [12] and its follow-ups [24, 13, 11]. A key finding in [12] establishes a convex tradeoff between perceptual quality and distortion in image restoration, applicable to any distortion measure and distribution. Moreover, perfect perceptual quality comes at the expense of no more than 3dB in PSNR. The work in [24] extends this, providing closed-form expressions for the tradeoff when MSE distortion and Wasserstein-2 distance are considered as distortion and perception measures respectively. In [51], it is shown that the Lipschitz constant of any deterministic estimator grows to infinity as it approaches perfect perception. Our work centers on uncertainty, reflecting the confidence level of a recovery algorithm. Thus, our works complements these studies by introducing uncertainty into the equation and exploring its impact on the perception-distortion relationship.
@@ 3 Problem Formulation
In this paper, we analyze estimators $\hat{X}$ with respect to two performance criteria: perception and uncertainty. To assess perceptual quality, we follow a theoretical approach, similar to previous works [74, 12], and measure perception using conditional divergence between $X$ and $\hat{X}$ defined as
@@
where $D_{v}$ stands for general divergence function. When an estimator attains a low value of the metric above, we say it exhibits high perceptual quality. When it comes to uncertainty, there are diverse practical methods to quantify it [29, 1]. However, for our analysis, we aim to identify a fundamental understanding of uncertainty. Therefore, we adopt the concept of entropy power from information theory, which assesses the statistical spread of a random variable. For the definition of entropy power and other relevant background, we refer the reader to Appendix A. Utilizing entropy power, we formally define the inherent uncertainty intrinsic to the restoration problem as follows
@@
Consider $Y=X+W$ where $X\sim{\mathcal{N}}(0,1)$ and $W\sim{\mathcal{N}}(0,\sigma^{2})$ are independent. Let the perception measure be the symmetric Kullback–Leibler (KL) divergence $D_{\text{SKL}}$ and assume stochastic estimators of the form $\hat{X}=\mathbb{E}\left[X|Y\right]+Z$ where $Z\sim{\mathcal{N}}(0,\sigma_{z}^{2})$ is independent of $Y$. As derived in Appendix B, the UP function admits a closed form expression in this case, given by
@@
The above result, illustrated in Appendix B, demonstrates the minimal attainable uncertainty increases as the perception quality improves. Moreover, The above example suggests a structure for uncertainty-perception function $U(P)$, which fundamentally relies on the inherent uncertainty $N(X|Y)$. Remarkably, the following section shows that this dependency generalizes beyond the specific example presented here, where its particular form is determined by the underlying distributions, along with the specific perception measure employed.
@@ 4.1 The Uncertainty-Perception Plane
The theorem establishes a fundamental tradeoff between perceptual quality and uncertainty in image restoration, regardless of the specific divergence measure, data distributions, or restoration model employed. This tradeoff is fundamentally linked to the inherent uncertainty $N(X|Y)$ arising from the information loss during the observation process. Notably, the upper bound of $N(X|Y)$ can be expressed as
@@
This shows that as $X$ approaches Gaussianity, $N(X|Y)$ approaches $N(X_{G}|Y)$. However, concurrently, it implies in general higher values of $N(X|Y)$ due to Lemma 1 of Appendix A. This finding yields a surprising insight: for multivariate Gaussian distributions, perfect perceptual quality comes at the expense of exactly twice the inherent uncertainty of the problem.
The next theorem shows that for a fixed perceptual index $P$, the optimal algorithms lie on the boundary of the constraint set. This facilitates the optimization task, as it restricts the search space to the boundary points.
@@
While we set $r=1/2$ to facilitate our derivations, it is important to note that all orders $r\in(0,1)$ are equivalent (see Appendix A). Consequently, given this equivalence and the close relationship between Rényi divergence and other metrics, analyzing the specific formulation provided by (5) may yield valuable insights applicable to a wide range of divergence measures. The following theorem provides lower and upper bounds for the UP function.
@@
Figure 2: The uncertainty-perception plane, derived from Theorem 3. The impossible region demonstrates the inherent tradeoff between perception and uncertainty, while other regions may guide practitioners toward estimators that balance perception and uncertainty, highlighting potential areas for improvement.
@@ 4.2 Revisiting the Distortion-Perception Tradeoff
4.3 Quantitative Demonstration
Empirical observations are the catalyst for our theoretical framework. Thus, we support our theoretical findings via quantitative evaluation of image super-resolution (SR) algorithms, following previous works [12, 11, 24, 51]. Utilizing the BSD100 dataset [50], the goal is predicting a high-resolution (HR) image from its low-resolution counterpart obtained through 4×4$\times$4 × bicubic downsampling. We evaluate various recent SR algorithms, including EDSR [45], ESRGAN [71], SinGAN [63], SANGAN [38], DIP [67], SRResNet/SRGAN variants [43], EnhanceNet [58], and Latent Diffusion Models (LDMs) with parameter $\beta\in[0,1]$ [57], where $\beta=0$ recovers DDIM [33] and $\beta=1$ recovers DDPM [65]. Our performance analysis focuses on uncertainty, and common measures of perception and distortion. Perception measures include projected distribution loss (PDL) [22], LPIPS [75], and Single-FID (SFID) – a patch-based variation of Fréchet inception distance (FID) [32] for image pairs where we randomly crop patches from each image and compute the standard FID between the groups. Distortion measures include MSE, peak signal-to-noise ratio (PSNR), and structural similarity index (SSIM) [72]. Estimating entropy in high dimensions poses significant challenges [42]. Hence, we utilize an upper bound for uncertainty, $N(\hat{X}_{G}-X_{G}|Y)$, discussed in Appendix E. This practical alternative simplifies the computation to calculating the geometric mean of the singular values of the error covariance. Additional results using direct entropy estimation are presented in Appendix G.
@@ 4.3 Quantitative Demonstration
Figure 4: Evaluation of SR algorithms. Top: Uncertainty-perception plane showing the tradeoff between perceptual quality and uncertainty (y-axis) for various perceptual measures. Bottom: Uncertainty-distortion plane illustrating the relation between uncertainty (x-axis) and distortion for different distortion measures. Axis placement differs in the two rows to highlight the distinct roles of uncertainty.
Figure 5: Visual comparison of algorithms on the uncertainty-perception plane. Algorithms are ordered from low to high uncertainty (left to right), with an accompanying increase in perceptual quality and distortion.
@@
Figure 4 reveals the results of our analysis. As observed in the top row, across various perceptual measures, an unattainable blank region exists in the lower left corner, indicating that no model simultaneously achieves both low uncertainty and high perceptual quality. Furthermore, an anti-correlation emerges near this region, where modest improvements in perceptual quality translate to dramatic increases in uncertainty. This observation suggests the existence of a tradeoff between uncertainty and perception. Additionally, the bottom row showcases a strong relationship between uncertainty and distortion across diverse measures, demonstrating that any increase in uncertainty leads to a significant rise in distortion. [^2] Note that MSE is a measure of distortion, whereas PSNR and SSIM are measures of inverse distortion; this accounts for the negative slope in the first two figures, and the positive slope in the third. Figure 5 visually reinforces these findings by presenting selected recoveries ordered from low to high uncertainty. The results show a clear increase in perceptual quality with increasing uncertainty, accompanied by a visually subtle increase in distortion.
@@ 5 Conclusion
This study established the uncertainty-perception tradeoff in generative restoration, demonstrating that high perceptual quality leads to increased uncertainty, particularly in high dimensions. We characterized this tradeoff and its fundamental relation to the inherent uncertainty of the problem, introducing the uncertainty-perception plane which may guide practitioners in understanding estimator performance. By extending our analysis to MSE distortion, we showed that the distortion-perception tradeoff emerges as a direct consequence of the uncertainty-perception tradeoff. Experimental results confirmed our theoretical findings, highlighting the importance of this tradeoff in image restoration.
@@ 6 Limitations
While our theoretical analysis is applicable to any signal restoration task with non-invertible degradation, we focused our empirical validation on image super-resolution for consistency with prior art. Therefore, expanding the experiments to include additional image-to-image tasks and domains such as audio, video, and text could reveal broader implications and applications of this tradeoff.
Our analysis is grounded in the theoretical framework of entropy as a measure of uncertainty. However, the practical estimation of high-dimensional entropy remains an active area of research. To address this, we utilize a tractable upper bound in Section 4.3. This alternative approach calls for further study, investigating its potential to effectively quantify uncertainty in analyzing the performance of existing algorithms. Furthermore, incorporating this upper bound into the design of new algorithms, whether as an objective function to minimize or a regularization term to constrain, could enable explicit control over the uncertainty-perception tradeoff, potentially leading to more reliable solutions.

### v3

@@ Abstract
The pursuit of high perceptual quality in image restoration has driven the development of revolutionary generative models, capable of producing results often visually indistinguishable from real data. However, as their perceptual quality continues to improve, these models also exhibit a growing tendency to generate hallucinations – realistic-looking details that do not exist in the ground truth images. Hallucinations in these models create uncertainty about their reliability, raising major concerns about their practical application. This paper investigates this phenomenon through the lens of information theory, revealing a fundamental tradeoff between uncertainty and perception. We rigorously analyze the relationship between these two factors, proving that the global minimal uncertainty in generative models grows in tandem with perception. In particular, we define the inherent uncertainty of the restoration problem and show that attaining perfect perceptual quality entails at least twice this uncertainty. Additionally, we establish a relation between distortion, uncertainty and perception, through which we prove the aforementioned uncertainly-perception tradeoff induces the well-known perception-distortion tradeoff. We demonstrate our theoretical findings through experiments with super-resolution and inpainting algorithms. This work uncovers fundamental limitations of generative models in achieving both high perceptual quality and reliable predictions for image restoration. Thus, we aim to raise awareness among practitioners about this inherent tradeoff, empowering them to make informed decisions and potentially prioritize safety over perceptual performance.
Figure 1: Illustration of Theorem 3. In restoration tasks, the minimal attainable uncertainty is lower bounded by a function that begins at the inherent uncertainty $U_{\text{Inherent}}$ of the problem (Definition 2) and graudally increases up to twice this value as the recovery approaches perfect perceptual quality.
@@ 1 Introduction
Restoration tasks and inverse problems impact many scientific and engineering disciplines, as well as healthcare, education, communication and art. Generative artificial intelligence [80, 38, 10] has transformed the field of inverse problems due to its unprecedented ability to infer missing information and restore corrupted data. In the realm of image restoration, the quest for high perceptual quality has led to a new generation of generative models, capable of producing outputs of remarkable realism, virtually indistinguishable from true images.
@@
Concerns surrounding hallucinations have prompted the development of uncertainty quantification methods, designed to evaluate the reliability of generated outputs. These approaches offer crucial insights into the model’s confidence in its predictions, empowering users to assess potential deviations from the original data and make informed decisions. Despite this progress, the relationship between achieving high perceptual quality and the extent of uncertainty remains an understudied area.
@@
    We establish a relationship between uncertainty and mean squared error (MSE) distortion, demonstrating that the uncertainty-perception trade-off induces the well-known distortion-perception trade-off [14] (Theorem 4).
  4. 4.
    We empirically validate all theoretical findings through experiments on image super-resolution and inpainting (Section 5), covering a broad spectrum of recovery algorithms, diverse metrics and data distributions. Our experimental results for image inpainting are illustrated in Figure 2.
Figure 2: Image inpainting results. Algorithms are ordered from low to high perception (left to right). Note the corresponding increased hallucinations and distortion. See Section 5 for details.
@@ 2 Related Work
Perception Quantification Perceptual quality in restoration tasks encompasses how humans perceive the output, considering visual fidelity, similarity to the original, and absence of artifacts. While traditional metrics like PSNR and SSIM [82] capture basic similarity, they miss finer details and higher-level structures. Learned metrics like LPIPS [87], VGG-loss [72], and DISTS [22] offer improvements but still operate on pixel or patch level, potentially overlooking holistic aspects. Recently, researchers have leveraged image-level embeddings from large vision models like DINO [17] and CLIP [62] to capture high-level similarity. Further advancements include HyperIQA [74] that leverages self-adaptive hyper networks to blindly assess image quality in the wild, while LIQE [88] and QAlign [84] utilize large language models to capture high-level semantic similarity and alignment between the restored and original images. Here, we follow previous works [58, 14, 31] and adopt a mathematical notion of perceptual quality defined as the divergence between probability densities.
@@
The Distortion-Perception Tradeoff The most relevant studies to our research are the work on the distortion-uncertainty tradeoff [14] and its follow-ups [23, 15, 13]. A key finding in [14] establishes a convex tradeoff between perceptual quality and distortion in image restoration, applicable to any distortion measure and distribution. Moreover, perfect perceptual quality comes at the expense of no more than 3dB in PSNR. The work in [23] extends this, providing closed-form expressions for the tradeoff when MSE distortion and Wasserstein-2 distance are considered as distortion and perception measures respectively. In [58], it is shown that the Lipschitz constant of any deterministic estimator grows to infinity as it approaches perfect perception.
This work uniquely emphasizes uncertainty in image restoration, distinguishing it from distortion. While distortion measures how close a restored image is to the original, uncertainty quantifies the confidence in the restoration itself. This distinction is crucial for decision-making, as high uncertainty can hinder informed choices, complementing existing research on perceptual quality and robustness.
@@ 3 Problem Formulation
In this paper, we analyze estimators $\hat{X}$ with respect to two performance criteria: perception and uncertainty. To assess perceptual quality, we follow a theoretical approach, similar to previous works [85, 14], and measure perception using conditional divergence [^1] See Appendix A for a brief explanation of how conditional divergence relates to human perception. between $X$ and $\hat{X}$ defined as
@@
where $D_{v}$ stands for general divergence function. When an estimator attains a low value of the metric above, we say it exhibits high perceptual quality. When it comes to uncertainty, there are diverse practical methods to quantify it [28, 1]. However, for our analysis, we aim to identify a fundamental understanding of uncertainty. Therefore, we adopt the concept of entropy power from information theory, which assesses the statistical spread of a random variable. For the definition of entropy power and other relevant background, we refer the reader to Appendix B. Utilizing entropy power, we formally define the inherent uncertainty intrinsic to the restoration problem as follows
@@
Consider $Y=X+W$ where $X\sim{\mathcal{N}}(0,1)$ and $W\sim{\mathcal{N}}(0,\sigma^{2})$ are independent. Let the perception measure be the symmetric Kullback–Leibler (KL) divergence $D_{\text{SKL}}$ and assume stochastic estimators of the form $\hat{X}=\mathbb{E}\left[X|Y\right]+Z$ where $Z\sim{\mathcal{N}}(0,\sigma_{z}^{2})$ is independent of $Y$. As derived in Appendix C, the UP function admits a closed form expression in this case, given by
@@
The above result, illustrated in Appendix C, demonstrates the minimal attainable uncertainty increases as the perception quality improves. Moreover, The above example suggests a structure for uncertainty-perception function $U(P)$, which fundamentally relies on the inherent uncertainty $N(X|Y)$. Remarkably, the following section shows that this dependency generalizes beyond the specific example presented here, where its particular form is determined by the underlying distributions, along with the specific perception measure employed.
@@ 4.1 The Uncertainty-Perception Plane
The theorem establishes a fundamental tradeoff between perceptual quality and uncertainty in image restoration, regardless of the specific divergence measure, data distributions, or restoration model employed. This tradeoff is fundamentally linked to the inherent uncertainty $N(X|Y)$ arising from the information loss during the observation process. Notably, the upper bound can be expressed as
@@
This shows that as $X$ approaches Gaussianity, $N(X|Y)$ approaches $N(X_{G}|Y)$. However, concurrently, it implies in general higher values of $N(X|Y)$ due to Lemma 1 of Appendix B. This finding yields a surprising insight: for multivariate Gaussian distributions, perfect perceptual quality comes at the expense of exactly twice the inherent uncertainty of the problem.
Next, we show that for a fixed perceptual index $P$, the optimal algorithms lie on the boundary of the constraint set. This facilitates the optimization, as it restricts the search space to the boundary points.
@@
Note that the assumption of the convexity of $\mathcal{D}_{v}$ in its second argument is not a restrictive condition. In fact, most widely-used divergence functions, notably all $f$-divergences (such as KL divergence, total variation distance, Hellinger distance, and Chi-square divergence), exhibit this property.
@@
While we set $r=1/2$ to facilitate our derivations, it is important to note that all orders $r\in(0,1)$ are equivalent (see Appendix B). Consequently, given this equivalence and the close relationship between Rényi divergence and other metrics, analyzing the specific formulation provided by (5) may yield valuable insights applicable to a wide range of divergence measures. The following theorem provides lower and upper bounds for the UP function.
@@
Figure 3: The uncertainty-perception plane (Theorem 3). The impossible region demonstrates the inherent tradeoff between perception and uncertainty, while other regions may guide practitioners toward estimators that better balance the two factors, highlighting potential areas for improvement.
@@ 4.2 Revisiting the Distortion-Perception Tradeoff
5 Experiments
Setup. Our theoretical framework is grounded in empirical observations, leading us to validate our findings through experiments on common benchmark tasks: image super-resolution and inpainting. We analyze performance through the lens of uncertainty, alongside established measures of perceptual quality and distortion. To assess perceptual quality, we employ state-of-the-art metrics including HyperIQA [74], LIQE [88] and Q-ALIGN [84]. Distortion is evaluated using traditional measures: MSE, peak signal-to-noise ratio (PSNR), and structural similarity index (SSIM) [82]. Accurately estimating entropy in high-dimensional spaces presents significant challenges [46]; hence, we utilize an upper bound for uncertainty, $N(\hat{X}_{G}-X_{G}|Y)$, as detailed in Appendix F. This practical alternative simplifies computation to calculating the geometric mean of the singular values of the error covariance.
For super-resolution, we utilize the BSD100 benchmark dataset [55], aiming to predict a high-resolution image from its low-resolution counterpart obtained via 4×4$\times$4 × bicubic downsampling. Our evaluation spans a diverse range of recovery algorithms, including EDSR [50], ESRGAN [81], SinGAN [71], SANGAN [39], DIP [77], SRResNet/SRGAN variants [47], EnhanceNet [66], and Latent Diffusion Models (LDMs) with parameter $\beta\in[0,1]$ [65], where $\beta=0$ recovers DDIM [32] and $\beta=1$ recovers DDPM [73]. In the context of image inpainting, we leverage the SeeTrue dataset [86], an image-text alignment benchmark known for its diverse collection of real and synthetic text-image pairs. Here, we focus our analysis on diffusion models due to their state-of-the-art performance and growing popularity in the field.
@@ 5 Experiments
Figure 5: Evaluation of SR algorithms. Top: Uncertainty-perception plane showing the tradeoff between perceptual quality and uncertainty (y-axis) for various perceptual measures. Bottom: Uncertainty-distortion plane showing the relationship between uncertainty and various distortion measures. Axis placement differs in the two rows to highlight the distinct roles of uncertainty.
@@
Results. Figure 5 presents our super-resolution analysis. As observed in the top row, across various perceptual measures, an unattainable blank region exists in the lower right corner, indicating that no model simultaneously achieves both low uncertainty and high perceptual quality. Furthermore, an anti-correlation emerges near this region, where modest improvements in perceptual quality translate to dramatic increases in uncertainty. This observation suggests the existence of a tradeoff between uncertainty and perception. Additionally, the bottom row showcases a strong relationship between uncertainty and distortion across diverse measures, demonstrating that any increase in uncertainty leads to a significant rise in distortion. [^3] Note that MSE is a measure of distortion, whereas PSNR and SSIM are measures of inverse distortion; this accounts for the negative slope in the first two figures, and the positive slope in the third. Figure 6 displays similar trends for image inpainting, consistent with our super-resolution analysis and reinforcing the validity of our findings across diverse restoration tasks and data distributions. This is further visualized in Figure 2, which presents outputs from selected algorithms ordered by perceptual quality. The results clearly demonstrate an increase in hallucination (uncertainty) and distortion with increasing perceptual quality. Finally, Appendix H presents additional results obtained via direct estimation of statistics in high dimensions, further supporting our theoretical analysis.
Figure 6: Evaluation of LDMs on image inpainting, highlighting the trade-off between uncertainty and perceptual quality (top) and the uncertainty-distortion relationship (bottom). No model achieves both low uncertainty and high perceptual quality, with higher uncertainty generally leading to increased distortion. Differing axis placements emphasize the distinct roles of uncertainty.
@@ 6 Conclusion
This study established the uncertainty-perception tradeoff in generative restoration, demonstrating that high perceptual quality leads to increased hallucination (uncertainty), particularly in high dimensions. We characterized this tradeoff and its fundamental relation to the inherent uncertainty of the problem, introducing the uncertainty-perception plane which may guide practitioners in understanding estimator performance. By extending our analysis to MSE distortion, we showed that the distortion-perception tradeoff emerges as a direct consequence of the uncertainty-perception tradeoff. Experimental results confirmed our theoretical findings, highlighting the importance of this tradeoff in image restoration.
@@ 7 Limitations
Our analysis is grounded in the theoretical framework of entropy as a measure of uncertainty. Information theory offers a powerful framework for quantifying uncertainty and dependencies in data, handling multivariate and heterogeneous data types, and capturing complex patterns. However, its wider adoption has been limited by the challenge of estimating information-theoretic measures in high dimensions. The curse of dimensionality makes accurate density estimation infeasible [12, 48], leading many to rely on simpler second-order statistics.
The development of practical tools for estimating statistics in high-dimensional data remains an active area of research [76]. While initial approaches assumed exponential family distributions (e.g., Gaussian) for tractable calculations [57], their performance degrades for long-tailed distributions. Non-parametric methods like binning strategies, including KDE and kNN estimators [61, 40, 29], offer more flexibility but are data-dependent and sensitive to parameter choices. Alternative approaches involve ensemble estimation [43] or von Mises Expansions [35], the distributional analog of the Taylor expansion. Rotation-Based Iterative Gaussianization [46] presents a promising direction by transforming data into a multivariate Gaussian domain, simplifying density estimation. However, its application to images has been limited to small patches due to the computational challenges of learning rotations based on principal or independent component analysis. A recent extension addresses this by utilizing convolutional rotations, enabling efficient processing of entire images [45].
While accurately estimating high-dimensional entropy remains an active research area, Section 5 utilizes a tractable upper bound. This alternative calls for further investigation into its potential for quantifying uncertainty and analyzing algorithm performance. Moreover, incorporating this bound into the design of new algorithms could enable explicit control over the uncertainty-perception trade-off, potentially leading to more reliable solutions.
Lastly, we focused our empirical validation on image super-resolution and inpainting, two benchmark problems in image restoration. Our analysis, however, applies to any restoration task with non-invertible degradation. Hence, expanding the experiments to additional image-to-image tasks and domains such as audio, video, and text may reveal broader implications and applications of our work.